Operation Summary for November 25, 2009 - November 30, 2009 |
The Integrated Luminosity for the period from 11/23/09 to 11/30/09 was 46.57 pb-1. NuMI reported receiving 7.6E18 protons on target during this same period. Wednesday Nov. 25 The day shift began with Operations monitoring the Tevatron (TeV) store 7385 (88.07E30) and the Recycler stash (251.1E10), with the Antiproton Source stacking (23.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7385 at 2:15 PM. Operations put the TeV into HEP shot setup, but that ended with a quench of the B0 low beta magnets during the squeeze. Water technicians made up water to the MI LCW loop following a Recycler stash. During the evening shift, Operations established beam to the Meson MTest area around 4:30 PM for beamline tuning. The MI LCW system was put into stand-alone mode for the holiday weekend due to pressure regulation issues at CUB. Instructions were left for watching temperatures and resistivity in MI over the weekend. Operations put the TeV through a dry squeeze at 7:30 PM. The NuMI extraction kicker tripped on fluorinert level just prior to 9:00 PM. Operations called in Mechanical Support personnel to fill the NuMI coolant skid; Operations had beam reestablished to NuMI by 10:45 PM. Operations checked the TeV orbits and then put the TeV into shot setup 11:45 PM. Thursday Nov. 26 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (217.1E10), with Pbar stacking (19.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7391 at 2 AM with an AIL of 197E30. A dipole magnet power supply at TeV sector F12 tripped off at 4:45 AM prompting CDF and DZero to put the detectors into standby due to beam position problems. Operators reset the dipole power supply and normal beam positions resumed. The experiments had their detectors back on by 5:20 AM. During the day shift, Pelletron corrector read-back problems held off antiproton transfers to the Recycler from 9:30 AM to 1:30 PM. During the evening shift, Operations terminated store 7391 at 6:20 PM. D0 made an access and the operators changed out the vertical dipole regulator at TeV sector F12. Operators checked the TeV orbits and had the TeV in shot setup by 9 PM. Operations established store 7394 at 10:15 PM with an AIL of 276E30. A Linac quadrupole power supply (Q64) developed a noisy/varying output regulation problem. Operators reseated the regulator card, which reduced the noise on the output. Regulation problems persisted however. Friday Nov. 27 The midnight shift began with Operations monitoring store 7394 (176.45E30) and the stash (136.2E10), with Pbar stacking (18.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators investigated the Linac quadrupole problem. Operators rebooted the MI RF low-level VME crate following a watchdog trip of the high-level RF system. During the day shift, Operations sent beam to the MTest area all shift. Operations terminated store 7394 at 11:45 AM. Operators prepared CDF and D0 for an access. During the evening shift, Operations stopped sending beam to the MTest area at 6:03 PM. A problem with the Pbar Accumulator A3B7 magnet limited stacking from 8:30 to 9:36 PM. Operations began the TeV turn-on and checkout at 9:37 PM. Saturday Nov. 28 The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (453.7E10), with Pbar stacking (90.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations checked the orbits, put the TeV into shot setup, and established store 7399 at 1:46 AM with an AIL of 281.9E30. Operations began sending beam to the MTest area at 4:40 AM. During the day shift, at 2:27 PM, MIRF2 tripped off on a cavity LCW leak and would not reset; operators bypassed the station and contacted RF personnel. They came in and brought MIRF15 back online. During the evening shift, Operations disabled the MTest beamline at 4:11 PM. MI resumed sending beam by 4:20 PM. At 5:34 PM, the P1 transfer line suffered a vacuum burst. Operations terminated store 7399 at 8:28 PM. Operations put the TeV into shot setup and established store 7400 at 10:11 PM with an AIL of 290.5E30, but there were problems with the antiproton beam positions at CDF. Sunday Nov. 29 The midnight shift began with Operations monitoring store 7400 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 1:22 AM, Operations reported a problem with BRF9; operators contacted an RF person and with his help discovered and replaced a bad power supply. Operations began sending beam to the MTest area at 4:24 AM. At 7:58 AM, the MI Power Supply interlock Loop tripped off due to lightning; it reset. During the day shift, Operations terminated store 7400 at 10:56 AM. Operators conducted TeV proton studies and then put the TeV into shot setup at 2:06 PM. Operations established store 7404 at 3:36 PM with an AIL of 261.11E30. During the evening shift, Operations stopped sending beam to the MTest area at 7:59 PM. At 10:16 PM, LRF2 tripped off; it reset. Monday Nov. 30 The midnight shift began with Operations monitoring store 7404 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7404 at 3:05 AM. Operations prepared D0 for access. Operations began the TeV turn-on and checkout at 4:33 AM. Operations put the TeV into shot setup and established store 7407 at 7:14 AM with a D0 luminosity of 269.2E30. (CDF did not report a luminosity due to beam position problems.) The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow a MI slipstacking study. |
|
More Information Comments and Suggestions |
| last modified 12/01/2009 email Fermilab |
| Security, Privacy, Legal | |